# Impact of positrons on electrical conductivity of hot and dense astrophysical plasma

### v1

@@ IV.2 Electrical conductivity from electron-ion and positron-ion collisions
As seen from Fig. 5, the position of the minimum around $T\simeq T_{F}/3$ remains unchanged also when positrons are present. However, the slope of the curves above the transition point rises abruptly, which is the consequence of the fast opening of the phase space for the pair creation with increasing temperature, as seen also from the phase diagram in Fig. 3. In contrast to the previous scaling $\sigma\propto T^{0.8}$ in the nondegenerate regime, we find here $\sigma\propto T^{4}$. To give a quantitative example, we compare the results with and without positrons at the lowest density considered here, $\rho=10^{6}\,\mathrm{g\,cm^{-3}}$. The inclusion of positrons increases the conductivity rougly by two orders of magnitude at $T=1$ MeV and by six orders of magnitude at $T=10$ MeV.
IV.3 Electrical conductivity including electron-positron collisions
Figure 6: Same as Fig. 4, but with inclusion of epep collisions.
We now discuss the electrical conductivities obtained in the full calculation, where electron–positron collisions are included as well. Let us recall that the electron–positron and positron–electron collision rates given by Eqs. (75) and (86) contain two distinct microscopic interaction channels: ordinary scattering and pair annihilation and creation processes. The latter is found to contribute significantly less to the total collision rate than the pure scattering channel.
@@ IV.3 Electrical conductivity including electron-positron collisions
Figure 6 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$ and their sum $\sigma$ as functions of density for the same temperature values as in Fig. 4. Comparing these two figures, we see that the effect of electron-positron collisions becomes crucial in the low-density and high-temperature regime of non-degenerate matter where $T\gg T_{F}$, as expected. Here, the conductivity becomes density-independent. This behavior can be explained by the fact that $ep$ collisions heavily outweigh $ei$ and $pi$ collisions in this regime. As can be seen from Eq. (38), in this case the relaxation times contain only the collision rates $\nu_{ep}\simeq\nu_{pe}$, which are density independent in the regime $\mu_{e}\ll T$, thus resulting in density-independent conductivities in accordance with Eq. (91). In the degenerate regime, the electron-positron collisions are suppressed, and we recover the results of the previous section. Although the location of the minimum in the total conductivity remains unaffected by electron-positron collisions, these are non-negligible already at $T\geq T_{F}/3$.
Figure 7: Same as Fig. 5, but with inclusion of epep collisions.
@@
Turning to the temperature dependence of conductivities shown in Fig. 7, we see again that the main difference as compared to Fig. 5 appears in the high temperature regime above the minimum point $T\simeq T_{F}/3$. At very high temperatures, where $T\gg T_{F}$, the conductivity $\sigma$ becomes essentially independent of density. On dimensional grounds, one then finds the scaling $\sigma\propto T$ in this regime, which agrees with the high-$T$ asymptotics shown in Fig. 7. This behavior is particularly evident at $\rho=10^{6}~\mathrm{g\,cm^{-3}}$. Immediately above the transition temperature $T\simeq T_{F}/3$, $\sigma$ follows the power-law dependence $\sigma\propto T^{4}$ as the collisions with ions are still dominant here. However, this fast increase is mitigated at higher temperatures as a result of intense electron-positron collisions, which surpass the $ei$ and $pi$ collision rates. As a numerical example, we find that at $\rho=10^{6}$ g cm-3 the ratios of conductivities we find to those computed earlier in Ref. Harutyunyan2016 are factor $20$ and $30$ for $T=1$ and $T=10$ MeV, respectively.
@@
These factors decrease at higher densities but still reach nearly an order of magnitude across most of the outer crust at $T=10$ MeV. Concluding this section, we emphasize that the accounting for creation of positrons is essential for accurate determination of the transport properties in the outer crust of neutron stars, already at temperatures close to the Fermi temperature.
@@ V Summary
In this work, we computed the electrical conductivity of hot QED plasma relevant to the outer crusts of neutron stars. The temperatures and densities studied here cover the transition from the degenerate to the nondegenerate regime for electrons. The novelty of our work is the inclusion of the contribution of positrons to the conductivity, which becomes increasingly important as the temperature increases. In this regime, the conductivity is dominated by thermally created electron-positron pairs which scatter off the correlated nuclei as well as off each other via screened electromagnetic force.
Solving a system of coupled Boltzmann kinetic equations for electrons and positrons, we expressed the conductivity in terms of electron and positron relaxation times. These contain electron-ion, positron-ion and electron-positron collisions rates, which were computed by using the relevant leading-order scattering diagrams which include all screening factors of strongly correlated crustal plasma.
Numerical results were obtained for conductivities for crustal matter composed of $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}. We found that the conductivity increases with density in the degenerate regime, in agreement with earlier results of Ref. Harutyunyan2016 , but in the nondegenerate regime it shows a decreasing behavior, in contrast to the results of previous studies. As a result, the conductivity crosses a minimum at the transition point from the nondegenerate to the degenerate regime. As a function of temperature, the conductivity grows with a power-law $\sigma\propto T^{4}$ in the semi-degenerate regime close to the transition temperature $T\simeq T_{F}/3$, because of intense creation of thermal electron-positron pairs. In this region, the number of pairs increases as $\propto T^{3}$, but the scattering is still mainly due to ions. At higher temperatures $T\gg T_{F}$, the electron-positron collisions become more important than electron-ion and positron-ion collisions, resulting in a slower, almost linear increase of $\sigma$ with temperature. We find that the inclusion of positrons enhances the electrical conductivity in the hot regions of the outer crust by factors of about 10. The results obtained in this work therefore, underscore the importance of accounting for positrons in the transport properties of heated plasma in neutron-star crusts.
Acknowledgements
@@
Appendix A The electron-positron scattering amplitude
In this Appendix, we will provide the details of the calculation of the spin-averaged electron-positron scattering amplitude. From Eqs. (50)–(III.1) we have
@@ Appendix A The electron-positron scattering amplitude
$\frac{1}{2}\sum\limits_{s_{1}s_{3}}J_{\mu}J_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{3}}\Big[p_{1\mu}p_{3\nu}+p_{1\nu}p_{3\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2})\Big].$ (101)
@@
Substituting the relevant component of the currents into Eq. (97) we obtain for the $t$-channel terms
@@
We now modify the nominators using the substitutions ${\bm{p}}_{3}={\bm{p}}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}^{\prime}+{\bm{q}}$ and $\epsilon_{3}=\epsilon-\omega$, $\epsilon_{4}=\epsilon^{\prime}+\omega$ (with $p\equiv p_{1}$ and $p^{\prime}\equiv p_{2}$)
@@
Defining the thee angles between the vectors $\bm{q}$}, $\bm{p}$} and $\bm{p}$$^{\prime}$ (see Fig. 8)
@@
$\frac{1}{2}\sum\limits_{s_{1}s_{2}}\tilde{J}_{\mu}\tilde{J}_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{2}}\Big[p_{1\mu}p_{2\nu}+p_{1\nu}p_{2\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2})\Big].$ (109)
@@
We next show that the matrix element for $pe$ collisions is the same as the one for $ep$ collisions. To check this, it is sufficient to interchange electrons and positrons, which corresponds to the following choice of Dirac 4-currents instead of Eqs. (55)–(56)
@@
Starting with these currents, a computation analogous to the one above leads then to exactly the same Eqs. (101) and (102), as well as Eqs. (109) and (110). Therefore, also the final forms of the scattering and annihilation terms in the matrix element are the same.
@@ Appendix B Evaluating the electron-positron collision rate
Substituting the electron-positron scattering amplitude (93) in the collision rate (41) we obtain
@@
where we used Eq. (107). Similar expression follows from Eq. (112) for the primed angles. The integrals of $\cos\varphi$ and $\cos^{3}\varphi$ over the interval $0,2\pi]$ vanish, while the integral of $\cos^{2}\varphi$ gives $\pi$ and the terms that do not depend on $\varphi$ integrate to $2\pi$. Therefore,

### v2

@@ IV.3 Electrical conductivity from electron-ion and positron-ion collisions
As seen from Fig. 8, the position of the minimum around $T\simeq T^{*}$ also remains unchanged when positrons are present. However, the slope of the curves above the transition point rises abruptly, which is the consequence of the fast opening of the phase space for the pair creation with increasing temperature, as seen also from the phase diagram in Fig. 3. In contrast to the previous scaling $\sigma\propto T^{0.8}$ in the nondegenerate regime, here we find $\sigma\propto T^{4}$. To give a quantitative example, we compare the results with and without positrons at the lowest density considered here, $\rho=10^{6}\,\mathrm{g\,cm^{-3}}$. The inclusion of positrons increases the conductivity roughly by 2 orders of magnitude at $T=1$ MeV and by 6 orders of magnitude at $T=10$ MeV.
IV.4 Electrical conductivity including electron-positron collisions
@@
We now discuss the electrical conductivities obtained in the full calculation, where electron–positron collisions are included as well. Figure 9 shows the partial conductivities $\sigma^{-}$ and $\sigma^{+}$ and their sum $\sigma$ as functions of density for the same temperature values as in Fig. 7. Comparing these two figures, we see that the effect of electron-positron collisions becomes crucial in the low-density and high-temperature regime of nondegenerate matter where $T\gg T_{F}$, as expected. Here, the conductivity becomes density independent, as a consequence of density independence of relaxation times in this regime and the relation (91). In the degenerate regime, the electron-positron collisions are suppressed, and we recover the results of the previous section. Although the location of the minimum in the total conductivity remains unaffected by electron-positron collisions, these are non-negligible already at $T\geq T^{*}$.
Figure 11: Total conductivity as a function of density for three different ionic compositions and two values of the temperature indicated in the plot.
Figure 12: Total conductivity as a function of temperature for three different ionic compositions and three values of the density indicated in the plot.
@@
Turning to the temperature dependence of conductivities shown in Fig. 10, we see again that the main difference as compared to Fig. 8 appears in the high-temperature regime above the minimum point $T\simeq T^{*}$. At very high temperatures, where $T\gg T_{F}$, the conductivity $\sigma$ becomes essentially independent of density. On dimensional grounds, one then finds the scaling $\sigma\propto T$ in this regime, which agrees with the high-$T$ asymptotics shown in Fig. 10. This behavior is particularly evident at $\rho=10^{6}~\mathrm{g\,cm^{-3}}$. Immediately above the transition temperature $T\simeq T^{*}$, $\sigma$ follows the power-law dependence $\sigma\propto T^{4}$, as the collisions with ions are still dominant here. However, this fast increase is mitigated at higher temperatures as a result of intense electron-positron collisions, which surpass the $ei$ and $pi$ collision rates. As a numerical example, at $\rho=10^{6}$ g cm-3, the ratios of conductivities we find to those computed earlier in Ref. Harutyunyan2016 are factors of $20$ and $30$ for $T=1$ and $T=10$ MeV, respectively. These factors decrease at higher densities but still reach nearly an order of magnitude across most of the outer crust at $T=10$ MeV.
To give an estimate of how our results would change depending on the crust composition, we show the density and temperature dependence of the total conductivity $\sigma$ for a crust composed of $^{\makebox[3.98613pt][r]{$\scriptstyle 4$}}_{\makebox[3.98613pt][r]{$\scriptstyle$}}\mathrm{He}$}, $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$} and $^{\makebox[7.97226pt][r]{$\scriptstyle 26$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Fe}$} in Figs. 11 and 12, respectively. The qualitative behavior of the conductivity is similar for these three elements. In all cases, there is a minimum (both in the density and temperature dependence) at the point where $T\simeq T^{*}$. At densities above or temperatures below this point, the conductivity is dominated by electron-ion collisions and scales as $\sigma\propto Z^{-1}$, as follows from Eqs. (74) and (90) (recall that $n_{e}=Zn_{i}=Z\rho/Am_{n}$ is essentially independent of $Z$). In the high-temperature and low-density regime, the conductivity becomes independent of both density and nuclei species, approaching a temperature-dependent value $\sigma\,[{\rm s}]\simeq 3\times 10^{22}\,T\,[{\rm MeV}]$. Thus, the composition dependence of conductivity is less pronounced in the high-temperature regime where electron-positron collisions play the main role in relaxation processes.
@@
Concluding this section, we emphasize that accounting for creation of positrons is essential for the accurate determination of the transport properties in the outer crust of neutron stars, already at temperatures close to the Fermi temperature.
@@ V Summary
In this work, we computed the electrical conductivity of hot QED plasma relevant to the outer crusts of neutron stars. The temperatures and densities studied here cover the transition from the degenerate to the nondegenerate regime for electrons. The novelty of our work is the inclusion of the contribution of positrons to the conductivity, which becomes increasingly important as the temperature increases. In this regime, the conductivity is dominated by thermally created electron-positron pairs, which scatter off the correlated nuclei as well as off each other via screened electromagnetic force.
Solving a system of coupled Boltzmann kinetic equations for electrons and positrons, we expressed the conductivity in terms of electron and positron relaxation times. These contain electron-ion, positron-ion and electron-positron collision rates, which were computed by using the relevant leading-order scattering diagrams, which include all screening factors of strongly correlated crustal plasma.
Numerical results were obtained for conductivities mainly for crustal matter composed of $^{\makebox[7.97226pt][r]{$\scriptstyle 12$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{C}$}. We found that the conductivity increases with density in the degenerate regime, in agreement with earlier results of Ref. Harutyunyan2016 , but in the nondegenerate regime, it shows a decreasing behavior, in contrast to the results of previous studies. As a result, the conductivity crosses a minimum at the transition point from the nondegenerate to the degenerate regime. As a function of temperature, the conductivity grows with a power law $\sigma\propto T^{4}$ in the semidegenerate regime close to the transition temperature $T\simeq T_{F}/3$ because of intense creation of thermal electron-positron pairs. In this region, the number of pairs increases as $\propto T^{3}$, but the scattering is still mainly due to ions. At higher temperatures $T\gg T_{F}$, the electron-positron collisions become more important than electron-ion and positron-ion collisions, resulting in a slower, almost linear increase of $\sigma$ with temperature. We find that the inclusion of positrons enhances the electrical conductivity in the hot regions of the outer crust by factors of about 10. The results obtained in this work therefore underscore the importance of accounting for positrons in the transport properties of heated plasma in neutron-star crusts.
The present study can be extended by considering more realistic, temperature-dependent crustal compositions that include multiple nuclear species in statistical equilibrium, as well as light nuclear clusters. Such an approach would provide a more accurate description of the microphysics in neutron star crusts, where the distribution of nuclei can vary significantly with density and temperature. In addition, the formalism developed here may be applied to improve calculations of various transport coefficients, including thermal conductivity, viscosities, and thermoelectric coefficients, in hot neutron star matter. Incorporating the effects of positrons in these calculations is particularly important, as they can significantly modify screening and collision rates, thereby influencing the transport properties of the warm crust. These extensions would allow for a more comprehensive and realistic modeling of the thermal and transport behavior of the hot transient state of compact stars at subnuclear densities.
ACKNOWLEDGMENTS
@@
Appendix A Electron-positron scattering amplitude
In this appendix, we will provide the details of the calculation of the spin-averaged electron-positron scattering amplitude. From Eqs. (50)–(III.1), we have
@@ Appendix A Electron-positron scattering amplitude
$\frac{1}{2}\sum\limits_{s_{1}s_{3}}J_{\mu}J_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{3}}\Big[p_{1\mu}p_{3\nu}+p_{1\nu}p_{3\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{3}-{\bm{p}}_{1}\cdot{\bm{p}}_{3}-m^{2})\Big],$ (101)
@@
Substituting the relevant components of the currents into Eq. (97), we obtain for the $t$-channel terms
@@
We now modify the numerators using the substitutions ${\bm{p}}_{3}={\bm{p}}-{\bm{q}}$, ${\bm{p}}_{4}={\bm{p}}^{\prime}+{\bm{q}}$ and $\epsilon_{3}=\epsilon-\omega$, $\epsilon_{4}=\epsilon^{\prime}+\omega$ (with $p\equiv p_{1}$ and $p^{\prime}\equiv p_{2}$)
@@
Defining the three angles between the vectors $\bm{q}$}, $\bm{p}$}, and $\bm{p}$$^{\prime}$ by (see Fig. 13)
@@
$\frac{1}{2}\sum\limits_{s_{1}s_{2}}\tilde{J}_{\mu}\tilde{J}_{\nu}^{*}$ $=$ $\frac{e^{*2}}{2\epsilon_{1}\epsilon_{2}}\Big[p_{1\mu}p_{2\nu}+p_{1\nu}p_{2\mu}-g_{\mu\nu}(\epsilon_{1}\epsilon_{2}-{\bm{p}}_{1}\cdot{\bm{p}}_{2}+m^{2})\Big],$ (109)
@@
We next show that the matrix element for $pe$ collisions is the same as the one for $ep$ collisions. To check this, it is sufficient to interchange electrons and positrons, which corresponds to the following choice of Dirac 4-currents instead of Eqs. (55) and (56),
@@
Starting with these currents, a computation analogous to the one above leads to exactly the same Eqs. (101) and (102), as well as Eqs. (109) and (110). Therefore, the final forms of the scattering and annihilation terms in the matrix element are the same.
@@ Appendix B Evaluating the electron-positron collision rate
Substituting the electron-positron scattering amplitude (93) in the collision rate (41), we obtain
@@
where we used Eq. (107). A similar expression follows from Eq. (112) for the primed angles. The integrals of $\cos\varphi$ and $\cos^{3}\varphi$ over the interval $0,2\pi]$ vanish, while the integral of $\cos^{2}\varphi$ gives $\pi$ and the terms that do not depend on $\varphi$ integrate to $2\pi$. Therefore,